# NorwAI’s Large Language Models: Technical Report

### v1

@@ 2 Available Models
Running the models. The models can be run directly with common open-source frameworks such as Hugging Face Transformers, vLLM, or Ollama. A minimal workflow employs Hugging Face Transformers with device_map=autodevice\_map=``auto" to distribute layers across available devices, as illustrated in Section 3.4.1. Quantized GGUF builds (Q3-Q8) are also supported, reducing memory requirements at a small cost in precision. Beyond the standard Transformers interface, models can be deployed using high-throughput servers such as vLLM, local or lightweight API hosting via Ollama, or integrated into managed infrastructures such as the Telenor AI Factory and SINTEF LM Studio. Typical (order-of-magnitude) GPU memory requirements are as follows: the smallest NorGPT-369M model runs comfortably on a 4 GB GPU or a modern CPU; models around 3 billion parameters typically require 12-16 GB of GPU memory; 7B-class models (Mistral and Llama2) perform best on 24-32 GB GPUs such as the A100 or RTX 4090; the Magistral-24B reasoning model requires approximately 64-80 GB; and the Mixture-of-Experts model Mixtral-8×7B is designed for distributed multi-GPU setups with around 115 GB or more total memory.
@@ 3.2 Pre-training
Pretraining relies on a carefully curated Norwegian corpus. For our initial corpus (NorLLM_Corpus_V1) created in 2023, we filtered Norwegian texts from the mC4 and OSCAR web-crawled corpora and included non-copyrighted material from the Norwegian National Library (Nasjonalbiblioteket) (Kummervold et al., 2021). We also sourced high-quality news articles from Schibsted and collected tweets (January 2012 to December 2022) as well as Reddit posts (October 2017 to December 2022) via their respective APIs. To enhance robustness in downstream tasks, we incorporated Danish, Swedish, and German texts, along with a small portion of English, all sourced from the mC4 corpus. The total corpus size was 2525B tokens.
@@ 3.3.1 Instruction Tuning and Alignment
Fine-Tuning on High-Quality Instruction Data We first fine-tuned the full model weights on a curated instruction dataset [^6] https://huggingface.co/datasets/mimir-project/mimir-instruction, which contains nearly 5,000 instructions annotated by our research assistants. The dataset spans various aspects of Norwegian culture, daily life, history, and geography. Tasks within the dataset include summarization, question answering, single-choice and multiple-choice questions, and other formats. Representative examples from this dataset are provided in the Appendix. Fine-tuning on this dataset enables the models to better understand user intent, adhere to instructions, and reflect Norwegian cultural nuances.
@@
Parameter-Efficient Fine-Tuning with LoRA To enhance efficiency and scalability across diverse applications, we employed LoRA (Low-Rank Adaptation), a parameter-efficient fine-tuning technique that reduces computational cost while maintaining high performance. Instead of updating all model parameters during fine-tuning, LoRA introduces a pair of small, trainable low-rank matrices within each transformer layer—typically in the attention and feed-forward projection weights. During training, only these low-rank parameters are optimized, while the original pre-trained weights remain frozen. This approach drastically reduces the number of trainable parameters and allows efficient adaptation to new tasks without degrading the base model’s general capabilities.
@@ 4.2.5 Evaluation Results on Schibsted News summarization
A human evaluation was conducted by Schibsted on 51 news articles, each accompanied by three summaries: one written by a professional journalist and two machine-generated summaries—one by GPT-4 and one by our Mixture-of-Experts model, NorwAI-Mixtral-8x7B-Instruct.
@@ 5 Model Selection and Usage Recommendations
The NorLLM family comprises models of varying sizes, architectures, and post‑training objectives. Selecting the appropriate model depends on task complexity, computational resources, and interaction requirements. The following guidelines summarize recommended usage scenarios.
5.1 Lightweight and General‑Purpose Use
NorGPT‑369M - Recommended for educational use, rapid prototyping, or environments with limited GPU memory (< 8 GB). Suitable for tokenization studies, sentence completion, and basic language modeling.
NorGPT‑3B / NorLlama‑3B - Balanced between efficiency and accuracy; appropriate for embedded applications, chatbots, and small‑scale summarization or classification tasks. Both models are trained from scratch on Norwegian data.
@@ 5.2 Advanced Comprehension and Generation
NorwAI-Mistral-7B (base) - A strong dense model for advanced comprehension and fluent Norwegian text generation. It outperforms NorwAI-Llama2-7B on NorEval and NorBench metrics and serves as a reliable backbone for high-quality analysis and synthesis tasks.
NorwAI-Mixtral-8×7B (base) - Designed for long-context generation (up to 32k tokens) and reasoning-heavy synthesis. The mixture-of-experts architecture dynamically activates two experts per token, improving performance and output quality on diverse downstream tasks relative to the NorwGPT series.
@@
5.3 Instruction‑Following and Dialogue Applications
NorwAI‑Mistral‑7B‑Instruct - Recommended as the default model for most Norwegian interactive and instruction‑following tasks. It balances accuracy and latency, supports a 4k‑token context, and provides reliable performance across a wide range of practical use cases.
NorwAI‑Mixtral‑8×7B‑Instruct - Ideal for complex assistant‑style dialogue, reasoning‑intensive tasks, and enterprise virtual‑agent deployments.
NorwAI‑Magistral‑24B - Performaning best on reasoning-heavy applications, combining broad knowledge with extended context. Especially designed for multi‑step reasoning, and tasks requiring long‑context understanding (up to 40k tokens). Recommended for research when longer inputs (16-32k tokens) or more complex reasoning are needed.
@@ 5.4 Domain and Resource Considerations
Low‑resource environments: Quantized GGUF versions (Q4 or Q5) of Mistral‑7B‑Instruct are recommended for laptop or edge inference.
High‑throughput serving: Deploy Mixtral‑8×7B or Magistral‑24B with vLLM or the Telenor AI Factory for scalable inference.
Domain adaptation: Apply LoRA fine‑tuning on top of Mistral‑7B‑Instruct or Mixtral‑8×7B‑Instruct for legal, healthcare, or educational applications to reduce compute cost.
Minority‑language tasks: Prefer NorwAI‑Magistral‑24B, trained on NorLLM_Corpus_V3, which includes substantial Nynorsk and Sámi data. NorwAI‑Mistral‑7B and NorwAI‑Mixtral‑8×7B were trained on NorLLM_Corpus_V2, which provides limited coverage but still basic support for mixed Norwegian text.

### v2

@@ 2 Available Models
Running the models. The models can be run directly with common open-source frameworks such as Hugging Face Transformers, vLLM, or Ollama. A minimal workflow employs Hugging Face Transformers with device_map=autodevice\_map=``auto" to distribute layers across available devices, as illustrated in Section 3.4.1. Quantized GGUF builds (Q3-Q8) are also supported, reducing memory requirements at a small cost in precision. Beyond the standard Transformers interface, models can be deployed using high-throughput servers such as vLLM, local or lightweight API hosting via Ollama, or integrated into managed infrastructures such as the Telenor AI Factory and SINTEF LM Studio. Typical (order-of-magnitude) GPU memory requirements are as follows: the smallest NorGPT-369M model runs comfortably on a 4 GB GPU or a modern CPU; models around 3 billion parameters typically require 12-16 GB of GPU memory; 7B-class models (Mistral and Llama2) perform best on 24-32 GB GPUs such as the A100 or RTX 4090; the Magistral-24B reasoning model requires approximately 64-80 GB; and the Mixture-of-Experts model Mixtral-8x7B is designed for distributed multi-GPU setups with around 115 GB or more total memory.
@@ 3.2 Pre-training
Pretraining relies on a carefully curated Norwegian corpus. For our initial corpus (NorLLM_Corpus_V1) created in 2023, we filtered Norwegian texts from the mC4 and OSCAR web-crawled corpora and included non-copyrighted material from the National Library of Norway (Nasjonalbiblioteket) [7]. We also sourced high-quality news articles from Schibsted and collected tweets (January 2012 to December 2022) as well as Reddit posts (October 2017 to December 2022) via their respective APIs. To enhance robustness in downstream tasks, we incorporated Danish, Swedish, and German texts, along with a small portion of English, all sourced from the mC4 corpus. The total corpus size was 2525B tokens.
@@ 3.3.1 Instruction Tuning and Alignment
Fine-Tuning on High-Quality Instruction Data We first fine-tuned the full model weights on a curated instruction dataset [^6] https://huggingface.co/datasets/mimir-project/mimir-instruction, which contains nearly 5000 instructions annotated by our research assistants. The dataset spans various aspects of Norwegian culture, daily life, history, and geography. Tasks within the dataset include summarization, question answering, single-choice and multiple-choice questions, and other formats. Representative examples from this dataset are provided in the Appendix. Fine-tuning on this dataset enables the models to better understand user intent, adhere to instructions, and reflect Norwegian cultural nuances.
@@
Parameter-Efficient Fine-Tuning with LoRA To enhance efficiency and scalability across diverse applications, we employed LoRA (Low-Rank Adaptation), a parameter-efficient fine-tuning technique that reduces computational cost while maintaining high performance. Instead of updating all model parameters during fine-tuning, LoRA introduces a pair of small, trainable low-rank matrices within each transformer layer, typically in the attention and feed-forward projection weights. During training, only these low-rank parameters are optimized, while the original pre-trained weights remain frozen. This approach drastically reduces the number of trainable parameters and allows efficient adaptation to new tasks without degrading the base model’s general capabilities.
@@ 4.2.5 Evaluation Results on Schibsted News summarization
A human evaluation was conducted by Schibsted on 51 news articles, each accompanied by three summaries: one written by a professional journalist and two machine-generated summaries - one by GPT-4 and one by our Mixture-of-Experts model, NorwAI-Mixtral-8x7B-Instruct.
@@ 5 Model Selection and Usage Recommendations
The NorLLM family comprises models of varying sizes, architectures, and post-training objectives. Selecting the appropriate model depends on task complexity, computational resources, and interaction requirements. The following guidelines summarize recommended usage scenarios.
5.1 Lightweight and General-Purpose Use
NorGPT-369M - Recommended for educational use, rapid prototyping, or environments with limited GPU memory (< 8 GB). Suitable for tokenization studies, sentence completion, and basic language modeling.
NorGPT-3B / NorLlama-3B - Balanced between efficiency and accuracy; appropriate for embedded applications, chatbots, and small-scale summarization or classification tasks. Both models are trained from scratch on Norwegian data.
@@ 5.2 Advanced Comprehension and Generation
NorwAI-Mistral-7B (base) - A strong dense model for advanced comprehension and fluent Norwegian text generation. It outperforms NorwAI-Llama2-7B on NorEval metrics and serves as a reliable backbone for high-quality analysis and synthesis tasks.
NorwAI-Mixtral-8x7B (base) - Designed for long-context generation (up to 32k tokens) and reasoning-heavy synthesis. The mixture-of-experts architecture dynamically activates two experts per token, improving performance and output quality on diverse downstream tasks relative to the NorwGPT series.
@@
5.3 Instruction-Following and Dialogue Applications
NorwAI-Mistral-7B-Instruct - Recommended as the default model for most Norwegian interactive and instruction-following tasks. It balances accuracy and latency, supports a 4k-token context, and provides reliable performance across a wide range of practical use cases.
NorwAI-Mixtral-8x7B-Instruct - Ideal for complex assistant-style dialogue, reasoning-intensive tasks, and enterprise virtual-agent deployments.
NorwAI-Magistral-24B - Performaning best on reasoning-heavy applications, combining broad knowledge with extended context. Especially designed for multi-step reasoning, and tasks requiring long-context understanding (up to 40k tokens). Recommended for research when longer inputs (16-32k tokens) or more complex reasoning are needed.
@@ 5.4 Domain and Resource Considerations
Low-resource environments: Quantized GGUF versions (Q4 or Q5) of Mistral-7B-Instruct are recommended for laptop or edge inference.
High-throughput serving: Deploy Mixtral-8x7B or Magistral-24B with vLLM or the Telenor AI Factory for scalable inference.
Domain adaptation: Apply LoRA fine-tuning on top of Mistral-7B-Instruct or Mixtral-8x7B-Instruct for legal, healthcare, or educational applications to reduce compute cost.
Minority-language tasks: Prefer NorwAI-Magistral-24B, trained on NorLLM_Corpus_V3, which includes substantial Nynorsk and Sámi data. NorwAI-Mistral-7B and NorwAI-Mixtral-8x7B were trained on NorLLM_Corpus_V2, which provides limited coverage but still basic support for mixed Norwegian text.